

********************************************************************
UARS WEEKLY STATUS REPORT - 29 APR 2005
********************************************************************
***********
SOLSTICE
***********
This week SOLSTICE has been under daylight only operations due to low
power availability. Additionally, SOLSTICE was powered of for about two
days because of concerns about red temperature limit excursions. Due to
SUSIM being off and our daylight only operations, the SOLSTICE
temperatures have dropped to a few degrees below the red limit.
Otherwise, the instrument is in good health gathering solar data on all
available orbital tracking periods. On average, we received 61% of each
day's data amounting to about 220 minutes of solar observing time per
day.
***********
SUSIM
***********
SUSIM was turned off on April 16 and remains off.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004366 has been reprocessed and archived at the GSFC DAAC. Additional
PEM telemetry is currently being processed.
MEPS data files are version 1. Processed MEPS data files at the DAAC
archive are between 1991274 and 2004366. Additional PEM telemetry is currently
being processed.
PEM continues to receive NMC data files from NOAA. NMC host machine
changed during mid-January which caused some NMC data transfer failures;
however, this situation was corrected and recovery completed with the
assistance of Roger Lin at NOAA. No problems have been encountered since
updates have occurred.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
PEM has received a request for data from the super storm of November
2004. The Dst value showed sudden commencement early on November 7, 2004
reaching a Dst value of about -380 gamma early on November 8, 2004.
PEM was not active at that time and became active late on November 8, 2004
in day only mode. A second peak of about -280 gamma was observed on November
11, 2004. PEM became active in day-light only mode late on November 8. PEM
survey spectrograms for November 8-18 of 2004 may be found at:
http://climatology.space.swri.edu/UARS_images_fromRudy.html
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 16 April 2005 (Orbit 74414, GMT Day 106) through
22 April 2005 (Orbit 74519, GMT Day 112).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
SUSIM 106/1523 74424 POWER OFF
PEM 107/0217 74430 ZEPS/NEPS OFF
SOLSTICE 107/0217 74430 FULLTIME TO DAY ONLY
HRDI 111/1617 74499 POWER OFF
The beta angle ranged between the angles of 51.6 degrees to 26.0
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 1 degree on 29 April 2005 (DOY 119).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 4.48 to 3.82 degrees C. Additional
battery performance data for each battery on 15 April 2005 and 22
April 2005 are provided in Appendix D.
The clock error ranged from 10.2 to -11.2 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 04
SUSIM = 01
PEM = 01
SOLSTICE = 01
HRDI = 05
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 03
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT made it through the first day of the HALOE turn on,
but then on day 106 the max differential on battery 2 rose to 308 mV.
At that time it was decided to power off SUSIM. This was executed on
the 106/1523 support. The next orbit following the SUSIM turn off,
the max differential on battery 2 dropped down to 190 mV. The
batteries remained fairly stable for the rest of the week until the
112/0310 event when the battery 2 max differential climbed to 280 mV
and fluctuated between 244-290 mV the remainder of the day.
Due to the high differential on battery 2, it was decided that day to
implement a new survival mode for HRDI to decrease their standby
power. The DAP was submitted and scheduled to take effect on 24 April
2005. Their new mode requires only the A side of their quiet bus to
be turned on for 30 minutes centered around midday. This is expected
to yield roughly an additional 35-40 watts, which should allow for
more margin to operate HALOE and SUSIM. The final results will be
seen next week.
Also this week, the FOT continued its testing of the solar array
drive. Once again both sides of the drive were utilized in order to
characterize their performance. The drive was moved 25 degrees in
either direction from 270. The results from these test runs were
inconsistent until the final days testing on 22 April 2005. The B
drive moved well from 266-294.4 without any signs of slipping or
stalling. This was the first time that the B drive operated this well
in the forward direction. It is too early to tell, but this could
mean that the A side has finally disengaged properly and in no longer
impeding the motion on the B drive. More testing will be done next
week to confirm this hypothesis.
MEETINGS
The FOT attended a teleconference on Friday April 22nd with the HRDI
PI's to discuss different options for their new survival mode of
operations.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 29 April 2005 (DOY 119). This
maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 22 April 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
74423 106/1455 05-131 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
74424 106/1523 05-132 SUSIM SUSIM POWER OFF
74430 107/0217 05-133 PEM ZEPS/NEPS OFF
74430 107/0217 05-134 SOLSTICE FULLTIME TO DAY-ONLY
74443 107/2211 05-135 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
74472 109/2023 05-136 SADDE SADDE TESTING
74499 111/1617 05-137 HRDI POWER HRDI OFF
74500 111/1655 05-138 HRDI PMON/RTS
74501 111/1756 05-139 HRDI PMON/RTS
74502 111/2016 05-140 HRDI HRDI PMON
74505 112/0056 05-141 HRDI HRDI HEATER OFF RTS-60
74514 112/1613 05-142 SADDE SADDE TESTING
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 22 April 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 22 April 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
74423 106/1455 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
74443 107/2211 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 22 April 2005
15 April 2005 - GMT Day 105 Beta = 52.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.94 +00.0/+00.0 06.9 0.00 0.0
2 291.2/ 44.8 +7.74/+5.79 +25.6/-17.2 25.8 1.04 19.2
3 190.4/-700.0 +2.91/+1.97 +20.0/-13.2 25.8 1.02 15.4
22 April 2005 - GMT Day 112 Beta = 26.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -112.0/-156.8 -3.68/-4.94 +00.0/+00.0 05.9 0.00 0.0
2 291.2/ -28.0 +7.08/+5.79 +31.6/-20.0 25.3 1.03 22.95
3 257.6/-700.0 +2.60/+1.97 +20.8/-12.0 25.1 1.04 14.10
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 22 April 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-023 FOT 74444 108/0035 30832 Generic Late Acquisition
05-023 FOT 74449 108/0748 30832 Generic Late Acquisition
05-025 FOT 74518 112/2157 39315 Intermittent Dropouts
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-023 FOT See List Apr 05 30832 Generic Late Acquisition
05-025 FOT 74518 112/2157 39315 Intermittent Dropouts